# Finding Transformer Circuits with Edge Pruning

### v1

@@ 1 Introduction
Mechanistic interpretability strives to understand models via bottom-up descriptions of their components (e.g., attention heads and MLPs in Transformers (Vaswani et al., 2017)). This typically proceeds via the identification and analysis of a circuit (Olah et al., 2020; Elhage et al., 2021)—a sparse computational subgraph of the model that captures the aspects of its behavior we wish to study. The arduous process of identifying circuits was recently automated by ACDC (Conmy et al., 2023) and EAP (Syed et al., 2023). However, ACDC uses an expensive greedy search that ablates each edge to estimate its importance. It cannot scale to datasets beyond a few hundred examples or to billion-parameter models. EAP, on the other hand, uses gradient-based linear approximations of activation patching to estimate the importance of all edges simultaneously. While fast, these approximations often sacrifice faithfulness to the full model. Besides, this approach ignores the impact of the presence/absence of other edges on the score.
@@ 2 Background: Circuit Discovery
The goal of circuit discovery is to facilitate a mechanistic understanding of Transformers by identifying the subset of a model’s computations that is most relevant to a particular model behavior. In this section, we define the computational graph of a Transformer, formalize the objective for circuit discovery and discuss the approaches of previous work.
@@
    ACDC  (Conmy et al., 2023) proposes to solve the above objective using greedy search—at each iteration, ACDC evaluates the effect of removing each edge individually, and removes any edge whose effect on the target metric is less than a specified threshold. This fails to capture the relative importance of similar edges and their interaction. Furthermore, the number of steps of the algorithm scales linearly with the number of edges, which is prohibitive at larger model sizes (e.g., CodeLlama-13B with $3.88$M edges).
@@
    Edge Attribution Patching (EAP)  (Syed et al., 2023) makes a linear approximation of activation patching to assign an importance score to each edge. This defines a ranking over edges, from which the top-$k$ edges are used to form a circuit of a specific sparsity. While the linear approximation can compute the importance scores efficiently in a single step, it is likely to find suboptimal solutions to the circuit discovery problem.
@@ 3 Method: Edge Pruning
We propose Edge Pruning, where we define masks not over nodes but over the edges connecting them. Specifically, we freeze the original model weights and introduce new trainable parameters $\boldsymbol{z}\in[0,1]^{|\mathcal{G}|}$, where $|\mathcal{G}|$ is the number of edges in the Transformer, and the parameter $z_{ji}$ is a relaxed binary mask for the edge $j\to i$. In other words, the pruning mask is an equivalent description of whether an edge is included ($z_{ji}=1$) or removed ($z_{ji}=0$) from the computational graph of a circuit. This formulation allows us to find subgraphs with greater granularity and precision compared to structured pruning, as the number of edges scales quadratically with the number of nodes in the computational graph of a model.
While structured pruning aims to discard pruned nodes as wasteful computation, the application to interpretability calls for more careful treatment of missing nodes and edges. Specifically, the activation of a removed edge $j\to i$ should be replaced by the interchange activation obtained from the corrupted version of the example (Section 2). To allow gradient-based optimization, we model the process as the masks continuously interpolating between the clean and corrupted activation. Specifically, we parameterize the $i$’th component as,
$y_{i}=f_{i}\left(z_{0i}y_{0}+(1-z_{0i})\tilde{y}_{0}+\sum_{j=1}^{i-1}\left(z_{ji}y_{j}+(1-z_{ji}\right)\tilde{y}_{j})\right),$ (3)
@@
In practice, concatenation increases the GPU memory footprint during training compared to regular structured pruning (Appendix B), but it is necessary for optimizing over edges between nodes that are separated by many layers. Despite the memory overhead, we demonstrate in Section 5 that we can scale our method to large models by parallelizing training over multiple GPUs.
We directly optimize the objective in (2) by performing stochastic gradient descent with respect to the edge weights $\boldsymbol{z}$. The target sparsity is enforced via $L_{0}$ regularization with a Lagrangian term $\mathcal{L}_{0}(\boldsymbol{z},c)$. We leverage the formulation of Louizos et al. (2018) to model the masks as hard concrete parameters and to circumvent the non-differentiability of the L0 term. At the end of the training, the edge weights are converted to binary masks based on a threshold (e.g., $0.5$), which uniquely determines the produced circuit (Figure 1c). Appendix A provides further details.
@@ Methods.
We compare Edge Pruning with a KL loss to ACDC and EAP in our experiments. Both are outlined in Section 2. We do not compare to other pruning-based methods, as Conmy et al. (2023) found them to perform much worse than ACDC. We list the hyperparameters used in Appendix B. The experiments in this section are all performed on GPT-2 Small (117M).
@@
A circuit is faithful to model behavior on a task if we can corrupt all model edges outside the circuit while retaining the model’s outputs (Hanna et al., 2024). We corrupt non-circuit edges with interchange ablation and evaluate the methods’ faithfulness as the KL divergence between model and circuit outputs. Specifically, we corrupt an example by swapping the placeholder value in the same template with a random example from the dataset. We appraise the circuits’ performance on IOI-t1, IOI, and GP via the Logit Difference $\log P(\text{correct})-\log P(\text{misleading})$ between the correct and misleading name/pronoun. For GT, we evaluate the Probability Difference $P(yy+1:99)-P(00:yy-1)$ between the correct and incorrect ranges. All metrics on GT work with predictions restricted to the set $\{00,01,\ldots,99\}$. We always take unrestricted predictions over the entire model vocabulary on other tasks. All non-Tracr experiments use a GPT-2 Small model. Appendix C evaluates the circuits on additional metrics.
@@ 4.2 Results
This section compares the three methods on our primary faithfulness and performance metrics. We report additional metrics in Appendix C, and Appendix F shows some circuits found by Edge Pruning.
@@
Edge Pruning is competitive on IOI-t1 and GP in terms of faithfulness at low sparsities, and slightly better at higher sparsities (Figure 2). It is considerably more faithful on IOI and GT than both ACDC and EAP, especially at higher sparsities. In particular, ACDC does worse than randomly choosing between the two names (KL divergence $0.69$) at high sparsities on IOI, whereas Edge Pruning remains decent. We hypothesize that the relative simplicity of IOI-t1 and GP—one template or small output space (he/she)—renders local (ACDC) or linear (EAP) approximations good proxies, potentially explaining the edge of Edge Pruning on IOI and GT. A similar trend is seen in performance (Figure 3): Edge Pruning finds better-performing circuits on all four tasks. Specifically, on IOI, Edge Pruning finds a circuit of $98.8\%$ sparsity that is as faithful and performs as well as the one found by ACDC at $96.8\%$ sparsity—using over 2.65×2.65$\times$2.65 × fewer edges. Interestingly, EAP scales better to higher sparsities than ACDC on GT, delivering respectable performance even at $99.5\%$ sparsity.
@@
We investigate how the methods scale to more examples at representative sparsities. To this end, we create a large version of the IOI dataset’s train split with 100K examples. We hold the number of gradient descent steps for Edge Pruning fixed (Appendix B). Although its runtime would scale linearly with more epochs, at 100K examples all approaches see almost all examples once. [^4] With our hyperparameters, Edge Pruning sees $96$k unique examples (can be higher with more GD steps). Thus, the time reported in Table 1 represents the relative overhead of each method. ACDC shows clear improvements with more examples, but cannot scale well due to prohibitive runtime. EAP, on the other hand, is fast even with more examples. However, it underperforms the other two methods significantly. Edge Pruning efficiently uses more examples and demonstrates both the least runtime and the highest faithfulness by far with $100$k examples. We therefore conclude that Edge Pruning is a good fit for complex or mixture distributions where more examples may be needed to specify model behavior.
@@
To check if Edge Pruning can find the ground-truth circuits, we use Tracr (Lindner et al., 2023) to compile two example programs—xproportion and reverse—as Transformers. The former yields a 2-layer Transformer that outputs, at each position, the fraction of x’s seen so far. The latter yields a 3-layer Transformer that can reverse lists. We use zero ablation following Conmy et al. (2023) (more details in Appendix B). Edge Pruning achieves perfect reconstruction of both circuits (Figure 4).
@@ 5 Case Study: Scaling to 13B Parameters
We work with the task Boolean Expressions from the BBH (Suzgun et al., 2022) benchmark suite. This task consists of instances of the form “((not False) and False) or (False and True) is $\rightarrow$ True”. The original dataset only has $250$ examples, so we programmatically generate an in-house version of the task. Our dataset has $3840,767$, and $3070$ examples in the train, validation, and test splits respectively. Each instance has between $3$ and $6$ literals, with a maximum nesting depth of $3$ and at most $2$ consecutive nots. We use $3$ demonstrations for the few-shot setting. The prompts used for the instruction-prompted and few-shot settings are provided in Appendix E. Our model is the instruction-finetuned version of CodeLlama-13B. [^5] https://huggingface.co/codellama/CodeLlama-13b-Instruct-hf It achieves accuracies of $82\%$ and $89.25\%$ in the instruction-prompted (IP) and few-shot (FS) settings, respectively.
@@
We next apply Edge Pruning to the described settings. We isolate one circuit when instruction prompting and one with the few-shot prompt (hyperparameters in Appendix B, which also highlights other optimizations like distributed training and gradient checkpointing). The circuit discovered in the IP setting has $1,041$ edges, corresponding to a $99.97\%$ edge sparsity. That discovered in the FS setting has $1,464$ edges, equivalent to $99.96\%$ edge sparsity. The discovered circuits are evaluated in Table 2. Despite using less than $0.04\%$ of the edges, the circuits closely match the performance of the full model—at $87.25\%$, the few-shot circuit performs within $2\%$ of the full model (when prompted few-shot). The instruction-prompted circuit is accurate within $2.75\%$ of the full model.
@@
We appraise the intersection of the IP and FS circuits next. It has $653$ edges, accounting for $62.7\%$ of the edges of the sparser (instruction prompted) circuit—this corresponds to an intersection over 1,700×1,700$\times$1 , 700 × larger than expected by random chance. We further evaluate the circuit formed by this intersection in the instruction prompted and few-shot settings (Table 2). It performs well in the instruction prompted setting, and worse than the model (but still significantly above chance) when prompted few-shot.
@@
Our case study reveals that the same mechanism (and the same circuit represented by the intersection above) explains a large part of the performance in both settings—i.e., they do not proceed via disjoint mechanisms. However, the performance gap between the FS and IP $\cap$ FS circuits is still sizable. Further, we see modest drops in cross-evaluation—e.g., from $87.25\%$ when evaluating the FS circuit few-shot to $75.75\%$ in the instruction prompted setting. This suggests that additional components are needed to complete the picture. A complete mechanistic description of the components in the two circuits is an exciting avenue for future work, but beyond the scope of this case study.
@@ 6 Related Work
By reducing a large model to a sparse subgraph, circuits help interpret internal model computations  (Olah et al., 2020; Elhage et al., 2021), and several visualization tools have been developed to aid this process (Sakarvadia et al., 2023; Katz and Belinkov, 2023; Tufanov et al., 2024). Circuits were originally found manually (Hanna et al., 2023; Athwin et al., 2023), but this has recently been automated by tools like ACDC (Conmy et al., 2023). ACDC uses activation patching (Vig et al., 2020) to knock out unimportant edges. Other approaches instead estimate the importance of each edge via attribution scores (Nanda, 2022) and was used by EAP (Syed et al., 2023). Ferrando and Voita (2024) use attribution patching to identify domain-specific model components in Llama-2-7B. Kramár et al. (2024) note that attribution patching may lead to incorrect approximations, and propose a variant with reduced error. In concurrent work, Hanna et al. (2024) argue that faithfulness metrics are better for evaluating circuits than measuring overlap with manual circuits. Recent work has explored other notions of a circuit. Inspired by the fact that Sparse Autoencoders (SAEs) can find human-interpretable features in LM activations (Cunningham et al., 2023)Marks et al. (2024) find circuits over these features.  Wu et al. (2023) align computation in Alpaca (Taori et al., 2023) with a proposed symbolic algorithm (Geiger et al., 2024). Our method is orthogonal to these developments.
@@
Pruning (LeCun et al., 1989) drops parameters or layers of a language model for space efficiency and potential speedups. Structured pruning (Wang et al., 2020; Xia et al., 2022) imposes some regularity on the resulting subnetworks, such as an equal fraction of preserved parameters in each layer. Doing so allows it to achieve substantial speedups on GPU hardware at the cost of lower compression. In contrast, unstructured pruning (LeCun et al., 1989; Hassibi and Stork, 1992) does not impose such constraints. Pruning has occasionally been used as part of an interpretability effort, but mostly at the level of neurons (Michel et al., 2019; Jain et al., 2023), or less commonly, attention heads/MLPs (Cao et al., 2021). Our work finds circuits by pruning the edges between components.

### v2

@@ 1 Introduction
Mechanistic interpretability strives to understand models via bottom-up descriptions of their components (e.g., attention heads and MLPs in Transformers (Vaswani et al., 2017)). This typically proceeds via the identification and analysis of a circuit (Olah et al., 2020; Elhage et al., 2021)—a sparse computational subgraph of the model that captures the aspects of its behavior we wish to study. The arduous process of identifying circuits (e.g., Wang et al. (2023)) was recently automated by ACDC (Conmy et al., 2023) and EAP (Syed et al., 2023). However, ACDC uses an expensive greedy search that ablates each edge to estimate its importance. It cannot scale to datasets beyond a few hundred examples or to billion-parameter models. EAP, on the other hand, uses gradient-based linear approximations of activation patching to estimate the importance of all edges simultaneously. While fast, these first-order approximations often sacrifice faithfulness to the full model. Besides, this approach ignores the impact of the presence/absence of other edges on the score.
@@ 2 Background: Circuit Discovery
The goal of circuit discovery is to facilitate a mechanistic understanding of Transformers by identifying the subset of a model’s computational graph that is most relevant to a particular model behavior. In this section, we define the computational graph of a Transformer, formalize the objective for circuit discovery, and discuss the approaches of previous work.
@@
    ACDC  (Conmy et al., 2023) proposes to solve the above objective using greedy search—at each iteration, ACDC evaluates the effect of removing each edge individually, and removes any edge whose effect on the target metric is less than a specified threshold. This fails to capture the relative importance of edges and their interaction. Furthermore, the number of steps of the algorithm scales linearly with the number of edges, which is prohibitive at larger model sizes (e.g., CodeLlama-13B with $3.88$M edges).
@@
    Edge Attribution Patching (EAP)  (Syed et al., 2023) makes a linear (first-order) approximation of activation patching to assign an importance score to each edge. This defines a ranking over edges, from which the top-$k$ edges are used to form a circuit of a specific sparsity. While the linear approximation can compute the importance scores efficiently in a single step, it is likely to find suboptimal solutions to the circuit discovery problem.
@@ 3 Method: Edge Pruning
We propose Edge Pruning, where we define masks not over nodes but over the edges connecting them. Specifically, we freeze the original model weights and introduce new trainable parameters $\boldsymbol{z}\in[0,1]^{|\mathcal{G}|}$, where $|\mathcal{G}|$ is the number of edges in the Transformer, and the parameter $z_{ji}$ is a relaxed binary mask for the edge $j\to i$. In other words, the pruning mask indicates whether an edge is included ($z_{ji}=1$) or removed ($z_{ji}=0$) from the computational graph of a circuit. This formulation allows us to find subgraphs with greater granularity and precision compared to structured pruning, as the number of edges scales quadratically with the number of nodes in a model’s computational graph.
While structured pruning discards pruned nodes by setting their activation to 0, the application to interpretability calls for more careful treatment of missing nodes and edges. Specifically, the activation of a removed edge $j\to i$ should be replaced by the interchange activation obtained from the corrupted version of the example (Section 2). To allow gradient-based optimization, we model the process as the masks continuously interpolating between the clean and corrupted activation. Specifically, we parameterize the $i$’th component as,
$y_{i}=f_{i}\left(z_{0i}y_{0}+(1-z_{0i})\tilde{y}_{0}+\sum_{\begin{subarray}{c}1\leq j<i\\ j\text{ upstream of }i\end{subarray}}\left(z_{ji}y_{j}+(1-z_{ji}\right)\tilde{y}_{j})\right),$ (3)
@@
In practice, concatenation increases the GPU memory footprint during training compared to regular structured pruning (Appendix A), but it is necessary for optimizing over edges between nodes that are separated by many layers. Despite the memory overhead, we demonstrate in Section 5 that we can scale our method to large models by parallelizing training over multiple GPUs.
We directly optimize the objective in (2) by performing stochastic gradient descent with respect to the edge weights $\boldsymbol{z}$. The target sparsity is enforced via $L_{0}$ regularization with a Lagrangian term. We leverage the formulation of Louizos et al. (2018) to model the masks as hard concrete parameters and to circumvent the non-differentiability of the L0 term. At the end of the training, the edge weights are converted to binary masks based on a threshold (e.g., $0.5$), which uniquely determines the produced circuit (Figure 1c). We now describe this process in more detail.
Details of the Edge Pruning process
Our formulation of pruning is based on that used by CoFi Pruning (Xia et al., 2022). Specifically, we model the masks $\boldsymbol{z}$ based on the hard concrete distribution as done by Louizos et al. (2018):
$\mathbf{u}\sim\text{Uniform}(\epsilon,1-\epsilon)$
$\tilde{\mathbf{s}}=\mathbf{s}\times(r-l)+l$
$\mathbf{z}=\min(1,\max(0,\tilde{\mathbf{s}}))$
Following, Wang et al. (2020), a target sparsity is enforced via a Lagrangian term (Louizos et al., 2018). If the current sparsity is $s$, the term, parametrized by a reference value $t$ is
$\mathcal{L}_{s}=\lambda_{1}\cdot(t-s)+\lambda_{2}\cdot(t-s)^{2}$
$\lambda_{1}$ and $\lambda_{2}$ are also updated during training via gradient ascent to keep the regularization tight. We vary the value of $t$ throughout training, linearly increasing it from 00 to a target value, as outlined in Appendix A. Although it may be useful to think of $t$ as a “target” sparsity, it is only a number. The runs usually converge to a value slightly below $t$, so it is prudent to set it to a value greater than $1$—although $s$ can then never reach the target value, it will be pushed to higher sparsities.
$\tilde{z}_{(n_{1},n_{2})}=z_{(n_{1},n_{2})}\times z_{n_{1}}$
We initially only used edge masks but found that the method would have difficulty converging to high sparsities (i.e., end up at low sparsities). Introducing a second set of masks allows the process to eliminate many edges quickly, accelerating the removal of unimportant components. However, the lagrangian above only applies to the edge masks. This is fine since the node masks can only remove further edges, not introduce new ones on top of those chosen by the edge masks. The final loss is
$\mathcal{L}=\mathcal{L}_{\text{KL}}+\mathcal{L}_{\text{edge},s}$
@@ Methods.
We compare Edge Pruning with a KL loss to ACDC and EAP in our experiments. Both are outlined in Section 2. We do not compare to other pruning-based methods, as Conmy et al. (2023) found them to perform much worse than ACDC. We list the hyperparameters used in Appendix A. The experiments in this section are all performed on GPT-2 Small (117M).
@@
A circuit is faithful to model behavior on a task if we can corrupt all model edges outside the circuit while retaining the model’s outputs (Hanna et al., 2024). We corrupt non-circuit edges with interchange ablation and evaluate the methods’ faithfulness as the KL divergence between model and circuit outputs. Specifically, we corrupt an example by swapping the placeholder value in the same template with a random example from the dataset. We appraise the circuits’ performance on IOI-t1, IOI, and GP via the Logit Difference $\log P(\text{correct})-\log P(\text{misleading})$ between the correct and misleading name/pronoun. For GT, we evaluate the Probability Difference $P(yy+1:99)-P(00:yy-1)$ between the correct and incorrect ranges. All metrics on GT work with predictions restricted to the set $\{00,01,\ldots,99\}$. We always take unrestricted predictions over the entire model vocabulary on other tasks. All non-Tracr experiments use a GPT-2 Small model. Appendix B evaluates additional metrics—including circuit overlap with manually found circuits.
@@ 4.2 Results
This section compares the three methods on our primary faithfulness and performance metrics. We report additional metrics in Appendix B, and Appendix F shows some circuits found by Edge Pruning.
@@
Edge Pruning is competitive on IOI-t1 and GP in terms of faithfulness at low sparsities, and slightly better at higher sparsities (Figure 2). It is considerably more faithful on IOI and GT than both ACDC and EAP, especially at higher sparsities. In particular, ACDC does worse than randomly choosing between the two names (KL divergence $0.69$) at high sparsities on IOI, whereas Edge Pruning remains better. We hypothesize that the relative simplicity of IOI-t1 and GP—one template or small output space (he/she)—renders local (ACDC) or first-order (EAP) approximations good proxies, potentially explaining the edge of Edge Pruning on IOI and GT. A similar trend is seen in performance (Figure 3): Edge Pruning finds better-performing circuits on all four tasks. Specifically, on IOI, Edge Pruning finds a circuit of $98.8\%$ sparsity that is as faithful and performs as well as the one found by ACDC at $96.8\%$ sparsity—using over 2.65×2.65$\times$2.65 × fewer edges. Interestingly, EAP scales better to higher sparsities than ACDC on GT, delivering respectable performance even at $99.5\%$ sparsity.
@@
We investigate how the methods scale to more examples at representative sparsities. To this end, we create a large version of the IOI dataset’s train split with 100K examples. We hold the number of gradient descent steps for Edge Pruning fixed (Appendix A). Although its runtime would scale linearly with more epochs, at 100K examples all approaches see almost all examples once. [^4] With our hyperparameters, Edge Pruning sees $96$k unique examples (can be higher with more GD steps). Thus, the time reported in Table 1 represents the relative overhead of each method. ACDC shows clear improvements with more examples, but cannot scale well due to prohibitive runtime. EAP, on the other hand, is fast even with more examples. However, it underperforms the other two methods significantly. Edge Pruning efficiently uses more examples and demonstrates both the least runtime and the highest faithfulness by far with $100$k examples. We therefore conclude that Edge Pruning is a good fit for complex or mixture distributions where more examples may be needed to specify model behavior.
@@
To check if Edge Pruning can find the ground-truth circuits, we use Tracr (Lindner et al., 2023) to compile two example programs—xproportion and reverse—as Transformers. The former yields a 2-layer Transformer that outputs, at each position, the fraction of x’s seen so far. The latter yields a 3-layer Transformer that can reverse lists. We use zero ablation following Conmy et al. (2023) (more details in Appendix A). Edge Pruning achieves perfect reconstruction of both circuits (Figure 4).
Edge Pruning is robust to variance in random initialization
Appendix D finds that both the resulting sparsity and the faithfulness of the circuits found by Edge Pruning are remarkably consistent across different random initializations of masks. We also investigate there the question of whether multiple different circuits can exist for a given task, and if Edge Pruning can find them.
@@ 5 Case Study: Scaling to 13B Parameters
We work with the task Boolean Expressions from the BBH (Suzgun et al., 2022) benchmark suite. This task consists of instances of the form “((not False) and False) or (False and True) is $\rightarrow$ False”. The original dataset only has $250$ examples, so we programmatically generate an in-house version of the task. Our dataset has $3840,767$, and $3070$ examples in the train, validation, and test splits respectively. Each instance has between $3$ and $6$ literals, with a maximum nesting depth of $3$ and at most $2$ consecutive nots. We use $3$ demonstrations for the few-shot setting. The prompts used for the instruction-prompted and few-shot settings are provided in Appendix E. Our model is the instruction-finetuned version of CodeLlama-13B. [^5] https://huggingface.co/codellama/CodeLlama-13b-Instruct-hf It achieves accuracies of $82\%$ and $89.25\%$ in the instruction-prompted (IP) and few-shot (FS) settings, respectively.
@@
We next apply Edge Pruning to the described settings. We isolate one circuit when instruction prompting and one with the few-shot prompt (hyperparameters in Appendix A, which also highlights other optimizations like distributed training and gradient checkpointing). The circuit discovered in the IP setting has $1,041$ edges, corresponding to a $99.97\%$ edge sparsity. That discovered in the FS setting has $1,464$ edges, equivalent to $99.96\%$ edge sparsity. The discovered circuits are evaluated in Table 2. Despite using less than $0.04\%$ of the edges, the circuits closely match the performance of the full model—the few-shot circuit achieves an accuracy of $87.25\%$ and performs within $2\%$ of the full model (when prompted few-shot). The instruction-prompted circuit is accurate within $2.75\%$ of the full model.
@@
We appraise the intersection of the IP and FS circuits next. The two circuits share $653$ edges, accounting for $62.7\%$ of the edges of the sparser (instruction prompted) circuit—this corresponds to an intersection over 1,700×1,700$\times$1 , 700 × larger than expected by random chance. We further evaluate the circuit formed by this intersection in the instruction prompted and few-shot settings (Table 2). It performs well in the instruction prompted setting, and worse than the model (but still significantly above chance) when prompted few-shot.
@@
Our case study suggests that the same mechanism (as represented by the intersection above) explains a large part of the performance in both settings—i.e., they do not proceed via disjoint mechanisms. However, the performance gap between the FS and IP $\cap$ FS circuits is still sizable. Further, we see modest drops in cross-evaluation—e.g., from $87.25\%$ when evaluating the FS circuit few-shot to $75.75\%$ in the instruction prompted setting. This suggests that additional components are needed to complete the picture. A complete mechanistic description of the components in the two circuits is an exciting avenue for future work, but beyond the scope of this case study.
Manual analysis of the CodeLlama-13B circuit.
Interpreting a circuit in such a large model—even if very sparse— remains a challenging task. We isolate a small region of the circuit and identify curious behavior in it in Appendix F, leading to an intriguing conjecture. Nonetheless, we believe that a thorough study requires more analysis, which is beyond the scope of this paper (but makes for exciting future work).
@@ 6 Related Work
By reducing a large model to a sparse subgraph, circuits help interpret internal model computations  (Olah et al., 2020; Elhage et al., 2021), and several visualization tools have been developed to aid this process (Sakarvadia et al., 2023; Katz and Belinkov, 2023; Tufanov et al., 2024). Circuits were originally found manually (Hanna et al., 2023; Athwin et al., 2023), but this has recently been automated by tools like ACDC (Conmy et al., 2023). ACDC uses activation patching (Vig et al., 2020) to knock out unimportant edges. Other approaches instead estimate the importance of each edge via attribution scores (Nanda, 2022); this approach was used by EAP (Syed et al., 2023). Ferrando and Voita (2024) use attribution patching to identify domain-specific model components in Llama-2-7B. Kramár et al. (2024) note that attribution patching may lead to incorrect approximations, and propose a variant with reduced error. In concurrent work, Hanna et al. (2024) argue that faithfulness metrics are better for evaluating circuits than measuring overlap with manual circuits. Recent work has explored other notions of a circuit. Inspired by the fact that Sparse Autoencoders (SAEs) can find human-interpretable features in LM activations (Cunningham et al., 2023)Marks et al. (2024) find circuits over these features.  Wu et al. (2023) align computation in Alpaca (Taori et al., 2023) with a proposed symbolic algorithm (Geiger et al., 2024). Our method is orthogonal to these developments.
@@
Pruning (LeCun et al., 1989) drops parameters or layers of a language model for space efficiency and potential speedups. Structured pruning (Wang et al., 2020; Xia et al., 2022) imposes some regularity on the resulting subnetworks, such as an equal fraction of preserved parameters in each layer. Doing so allows it to achieve substantial speedups on GPU hardware at the cost of lower compression. In contrast, unstructured pruning (LeCun et al., 1989; Hassibi and Stork, 1992) does not impose such constraints. Channel pruning (He et al., 2017) is a form of structured pruning that prunes input channels in vision models, which has been adapted for neural architecture search (e.g. Li et al., 2022). Pruning has occasionally been used as part of an interpretability effort, but mostly at the level of neurons (Michel et al., 2019; Jain et al., 2023), or less commonly, attention heads/MLPs (Cao et al., 2021). Our work finds circuits by pruning the edges between components instead.